# Feedback-Enhanced Online Multiple Testing with Applications to Conformal Selection

### v1

@@ 2.5 Addressing local dependence
We say that $p$-values $p_{1},p_{2},\dots,p_{t},\dots$ are local dependent if
@@
Zrnic et al. [2021] pioneered a strategy to handle this local dependency in online multiple testing, leading to the development of the $\text{LORD}_{\text{dep}}$ and $\text{SAFFRON}_{\text{dep}}$ methods. The following procedures are the modified counterparts of LF and SF under the local dependence:
$\widehat{\operatorname{FDP}}_{\text{LF}_{\text{dep}}}(t)$ $=\frac{\sum_{j\leq t-1}(1-\theta_{j})\alpha_{j}+\alpha_{t}}{(\sum_{j\leq t,j\notin\{t-L_{t},\dots,t-1\}}\delta_{j})\vee 1}\leq\alpha,$
$\widehat{\operatorname{FDP}}_{\text{SF}_{\text{dep}}}(t)$ $=\frac{\sum_{j<t-L_{t}}(1-\theta_{j})\alpha_{j}\frac{\mathbb{I}\{p_{j}>\lambda\}}{1-\lambda}+\sum_{j=t-L_{t}}^{t-1}(1-\theta_{j})\frac{\alpha_{j}}{1-\lambda}+\frac{\alpha_{t}}{1-\lambda}}{(\sum_{j\leq t,j\notin\{t-L_{t},\dots,t-1\}}\delta_{j})\vee 1}\leq\alpha.$
The corresponding testing levels are presented in Appendix C.1. The $\text{LF}_{\rm dep}$ and $\text{SF}_{\rm dep}$ methods control mFDR under local dependence.
@@
Suppose that the null $p$-values are locally dependent, as defined in (9). Then if the parameters $\{\alpha_{t}\}_{t\in\mathbb{N}}$ are selected such that $\widehat{\operatorname{FDP}}_{\operatorname{LF}_{\rm{dep}}}(t)\leq\alpha$ or $\widehat{\operatorname{FDP}}_{\operatorname{SF}_{\rm{dep}}}(t)\leq\alpha$, we have $\operatorname{mFDR}(t)\leq\alpha$ for all $t\in\mathbb{N}$.
3 Applications on Online Conformal Selection
The online conformal selection [Huo et al., 2024, Xu and Ramdas, 2024] provides a canonical and intuitive instantiation of the GAIF framework. Here, decisions must be made in real time about whether an incoming observation satisfies a pre-specified requirement, while the true label—i.e., the feedback—can generally be observed after the decision. Consider a data pair $({\bf X},Y)\in{\mathcal{X}}\times{\mathcal{Y}}$, with a historical calibration dataset $\mathcal{D}_{{\mathcal{C}}}=\{({\bf X}_{i},Y_{i})\}_{i=-n+1}^{0}$ of size $n$. Let $\mathcal{C}$} denote its index set. Test samples $({\bf X}_{t},Y_{t})$ arrive sequentially for $t=1,2,\dots$, where covariates ${\bf X}_{t}$ are observed but responses $Y_{t}$ remain hidden until a real-time decision $\delta_{t}\in\{0,1\}$ is made. The goal at each step is to determine whether $Y_{t}$ lies in a target region ${\mathcal{A}}\subseteq{\mathcal{Y}}$, (e.g., $[a,b]$ or $b,\infty)$ in regression tasks). This can be framed as an online testing problem with $\theta_{t}=\mathbb{I}\{Y_{t}\in{\mathcal{A}}\}$. We will specialize GAIF to this context: we first construct provably independent online conformal p-values, and then adapt GAIF rules to yield finite-sample mFDR control.
@@ 3.1 Construction of online conformal $p$-values
To construct $p$-values, we introduce the notion of non-conformity score function $V({\bf X})$, where larger values indicate a higher likelihood that $\theta_{i}=0$. Typically, $V({\bf X})$ is a monotone transformation of the prediction $\widehat{\mu}({\bf X})$, assumed pre-trained to to estimate $Y_{t}$. For example, if $\theta=\mathbb{I}\{Y_{t}\geq b\}$, then one can take $V(X)=b-\widehat{\mu}({\bf X})$. For simplicity, we write $V_{i}=V({\bf X}_{i})$. A natural approach to achieving online FDR control is to compute conformal $p$-values [Bates et al., 2023, Jin and Candès, 2023b] and then apply the GAIF procedure or existing GAI rules. However, this is not directly applicable in the present setting, as standard conformal $p$-values do not satisfy Eq.(8) or independence assumptions.
To circumvent these issues, we adopt online conformal $p$-values [Vovk et al., 2003, Vovk, 2021], which sequentially update the calibration set. Under exchangeability, this yields null $p$-values that are independent, thereby avoiding the complex dependence structures inherent in offline conformal methods, which rely on shared calibration sets and produce $p$-values with PRDS dependence [Bates et al., 2023]. We need the following assumption.
@@
This assumption is common in conformal inference [Marandon et al., 2024], which is weaker than requiring $\{({\bf X}_{i},Y_{i})\}_{i>-n}$ are independent and identically distributed (i.i.d.).
The strategy is that each test point is incorporated into the calibration set for the next time step. Specifically, at time $t$, the point $({\bf X}_{t},Y_{t})$ serves as the test point, but from time $t+1$ onward, it becomes part of the calibration data. Let ${\mathcal{C}}_{0t}$ denote the dynamically updated index set of calibration samples. Specifically, ${\mathcal{C}}_{0t}$ is given by
${\mathcal{C}}_{0t}=\{-n<i<t:\theta_{i}=0\}.$ (10)
where $\xi_{t}\overset{\text{i.i.d.}}{\sim}\operatorname{Unif}[0,1]$ is an independent random variable for tie-breaking.
Similar forms of online conformal $p$-values have been studied for testing exchangeability [Vovk et al., 2003, Vovk, 2021] and constructing online conformal prediction intervals [Angelopoulos et al., 2024]. We formally state the validity and independence under the null in Proposition 3.1.
@@
Moreover, we further allow the score function to vary over time and be data-dependent, provided it remains symmetric on null data; see Appendix A.3. This flexibility enables leveraging non-null feedback to improve score quality, facilitating model selection and adaptation to distribution shift; see Section 3.3.
3.2 Online conformal testing procedures
Having established the properties of online conformal $p$-values, we now introduce the corresponding testing procedures. Although applying GAIF directly to those $p$-values yields satisfactory FDR control empirically, theoretical guarantees remain lacking. The key challenge is that null online conformal $p$-values are still dependent with those under the alternative. To overcome this issue, we propose modified GAIF rules tailored for online conformal testing, offering rigorous, distribution-free, and model-agnostic mFDR guarantees.
The key idea is to construct the sequence $\{\alpha_{t}\}_{t\in\mathbb{N}}$ based solely on the rejections of true null hypotheses, rather than on all rejections. While this strategy may be conservative as the discard of non-null decision information, the integration of the feedback provides great compensation. Even when the proportion of non-nulls is relatively large, empirical evidence shows that our modified procedure can still achieve higher power than benchmarks that ignore feedback information.
We revise the rules in LF and SF by replacing $\tau_{j}$ with $\tilde{\tau}_{j}$, resulting in LFS and SFS rules, respectively, where $\tau_{j}$ denotes the time of the $j$-th rejection and $\tilde{\tau}_{j}$ denotes the time of the $j$-th rejection under the null, defined as
@@ 3.2 Online conformal testing procedures
The last letter S in “LFS” and “SFS” indicates “safe” since it guarantees rigorous finite-sample mFDR control. By doing so, $\tilde{\tau}_{j}$ is fixed given past null p-values, as its construction depends only on the rejections of true null hypotheses.
The Online Conformal Testing with Feedback (OCTF) procedure, based on online conformal $p$-values, is detailed in Algorithm 2. Building on the construction in Section 3.1 and the analysis above, we obtain finite-sample online mFDR control for the OCTF procedures, as formalized in Theorem 3.
Theorem 3 (Finite-sample Online mFDR control for OCTF).
Suppose Assumption 1 holds. The outputs $\mathcal{R}$} of Algorithm 2 satisfy $\operatorname{mFDR}(t)\leq\alpha$ for all $t\in\mathbb{N}$.
With respect to FDR control, the additional challenge lies in the fact that a past non-null decision $\delta_{i}=\mathbb{I}\{p_{i}\leq\alpha_{i}\}$ with $\theta_{i}=1$ may be dependent on a null $p_{t}$. FDR control requires decoupling the dependence between decisions $\sum_{i\leq t}\delta_{i}$ in the denominator and $p_{t}$, which is a more difficult task and warrants further research.
@@
1:  Initialize ${\mathcal{C}}_{01}=\{i\in{\mathcal{C}}:\theta_{i}=0\}$
@@
3:   Observe test data ${\bf X}_{t}$
4:   Compute conformity scores $\{V_{i}\}_{i\in{\mathcal{C}}_{0t}\cup\{t\}}$
5:   Compute online conformal $p$-value $p_{t}$ via (11)
6:   Update $\alpha_{t}={\alpha}_{t}^{\text{LFS}}$ in Eq. (25) (or $\alpha_{t}={\alpha}_{t}^{\text{SFS}}$ in Eq.(26))
7:   if $p_{t}\leq\alpha_{t}$ then $\delta_{t}=1$, else $\delta_{t}=0$
8:   Obtain the revealed feedback $Y_{t}$ (and thus obtain $\theta_{t}$)
@@
3.3 Optimized online conformal multiple testing with feedback
The performance of online conformal testing depends on the combined choice of the non-conformity scores $V({\bf X})$ and the prediction model (e.g., random forests, neural networks, support vector machines, or regularized linear models) [Bai and Jin, 2024, Gasparin and Ramdas, 2024]. We collectively refer to these choices as model selection. Moreover, in practice, data distributions in non-null may shift over time: the $\{({\bf X}_{t},Y_{t})\}_{t\in\mathbb{N}}$ remains exchangeable under the null but may change in distribution under the alternative [Gang et al., 2021, Huo et al., 2024]. Dynamically selecting models adapted to the current distribution can substantially alleviate the shifting effects. This motivates us to design a versatile framework that ensures error rate control in online conformal testing while allowing flexible data reuse for model optimization and ultimately enhancing power.
@@ 3.3 Optimized online conformal multiple testing with feedback
$\widehat{k}_{t}=\underset{k\in[K]}{\arg\min}\;{\mathcal{M}}(k,{\mathcal{D}}_{t}),$
where ${\mathcal{M}}(k,{\mathcal{D}}_{t})$ is a given criterion evaluated by the currently observed data ${\mathcal{D}}_{t}=\{({\bf X}_{i},Y_{i}):-n<i<t\}\cup\{{\bf X}_{t}\}$. And the $p$-value is generated by the optimized score function $V(\cdot;\widehat{k}_{t})$ for our online testing procedure. However, naively optimizing the predictive model can yield invalid online conformal $p$-values or introduce intricate dependencies, ultimately leading to FDR inflation [Zhang et al., 2022]. To ensure validity and independence of online conformal $p$-values after model-optimization, we propose a model evaluation criteria specifically designed for the online setting with feedback. In particular, our adaptive model selection approach accommodates distribution shifts in the non-null data by dynamically selecting the best model based on the recent alternative distribution.
Considering the availability of real-time feedback on past true signals, we propose using the exponentially weighted moving average (EWMA) of past auxiliary non-null $p$-values as the model selection criterion. At each time $t$, the optimal model is chosen by minimizing this EWMA value, which serves as a good proxy for the unknown current non-null $p$-value. The rationale of this criterion is that a better model should yield smaller non-null $p$-values, leading to more powerful procedures.
Specifically, at time $t$, for each score function $k$, we construct auxiliary non-null $p$-values as
$\tilde{p}^{k}_{j}=\frac{\sum_{s\in{\mathcal{C}}_{0t}\cup\{t\}}\mathbb{I}\{V(\mathbf{X}_{s};k)\leq V(\mathbf{X}_{j};k)\}}{1+|{\mathcal{C}}_{0t}|},\quad j\in{\mathcal{C}}_{1t},$ (12)
@@
where ${\mathcal{C}}_{1t}=\{-n<i<t:\theta_{i}=1\}$ denotes the set of online non-null samples observed prior to time $t$. The construction in (12) is carefully designed so that the resulting auxiliary non-null $p$-values are invariant under permutations of ${\mathcal{C}}_{0t}\cup\{t\}$, which is crucial for ensuring valid inference after model selection. We then define the model evaluation criterion as
${\mathcal{M}}^{\mathrm{EWMA}}_{t}(k,{\mathcal{D}}_{t})=\frac{\sum_{j\in{\mathcal{C}}_{1t}}\rho^{t-1-j}\,\tilde{p}^{k}_{j}\,}{\sum_{j\in{\mathcal{C}}_{1t}}\rho^{t-1-j}},$
@@
where $\rho\in(0,1)$ is the user-specified decay parameter that downweights past observations, enabling dynamic adaptation to the recent non-null distribution. The optimal model is accordingly selected as $\widehat{k}_{t}={\arg\min}_{k\in[K]}{\mathcal{M}}^{\mathrm{EWMA}}_{t}(k,{\mathcal{D}}_{t}).$
Finally, the online conformal $p$-value after model optimization, denoted as $p^{\rm opt}_{t}$, is computed by Eq.(11) with $V(\mathbf{X}_{t};\widehat{k}_{t})$ instead of $V_{t}$, and then the OCTF procedures in Section 3.2 can be implemented with these optimized online conformal $p$-values. The entire optimized online conformal testing procedure (Opt-OCTF) is summarized in Algorithm 3. Intuitively, this approach helps the $p$-value construction dynamically adapt as new data arrives, while ensuring that the past samples are still considered, but with diminishing influence. By continually learning the current non-null pattern, our procedure remains powerful against potential non-null distribution shifts.
Algorithm 3 Optimized Online Conformal Testing with Feedback (Opt-OCTF)
0:  Initial calibration data ${\mathcal{D}}_{\mathcal{C}}=\{({\bf X}_{i},Y_{i})\}_{i=-n+1}^{0}$, target region $\mathcal{A}$, FDR target level $\alpha\in(0,1)$, $K$ pretrained models $\{\widehat{\mu}_{k}\}_{k=1}^{K}$ and the non-conformity score function $V(\cdot;k)$, evaluating criterion $\mathcal{M}={\mathcal{M}}^{\text{EWMA}}$, parameter $s_{0}$, parameter sequence $\{\gamma_{t}\}_{t\in\mathbb{N}}$, constant parameters $\rho\in(0,1)$, $\lambda\in(0,1)$.
1:  Initialize ${\mathcal{C}}_{01}=\{i\in{\mathcal{C}}:\theta_{i}=0\}$
2:  for $t=1,2\ldots$ do
3:   Observe test data ${\bf X}_{t}$
4:   Decide the predictive model for $t$-th test sample by $\widehat{k}_{t}={\arg\min}_{k\in[K]}{\mathcal{M}}^{\text{EWMA}}(k,{\mathcal{D}}_{t})$
5:   Construct optimized online conformal p-value $p^{\rm opt}_{t}$
6:   Update $\alpha_{t}=\alpha_{t}^{\text{LFS}}$ in Eq.(25) (or $\alpha_{t}=\alpha_{t}^{\text{SFS}}$ in Eq.(26))
7:   Make a decision $\delta_{t}=\mathbb{I}\{p_{t}^{\text{opt}}\leq\alpha_{t}\}$
8:   Obtain the revealed feedback $Y_{t}$ (and thus obtain $\theta_{t}$)
9:   Update the calibration dataset ${\mathcal{C}}_{0t}$.
10:  end for
10:  Rejection set $\mathcal{R}_{\text{opt}}=\{t:\delta_{t}=1\}$.
Importantly, the optimized online conformal $p$-values still remain valid and mutually independent under the null hypothesis, which is an extension of Proposition 3.1 by allowing that the score function at each time to be dependent on the non-null data. As null $p$-values are excluded from the computation of EWMA, we can conclude the score function $V(\cdot;\hat{k}_{t})$ is symmetric to the null data up to time $t$. Consequently, Opt-OCTF still achieves finite-sample FDR control, as established in the following Corollary 3.1.

### v2

@@ 3.2 GAIF under local dependence
We say that $p$-values $p_{1},p_{2},\dots,p_{t},\dots$ are locally dependent if
@@
Zrnic et al. [2021] pioneered a strategy to handle this local dependence in online multiple testing, leading to the development of the $\text{LORD}_{\text{dep}}$ and $\text{SAFFRON}_{\text{dep}}$ methods. We extend the idea to the feedback-enhanced setting. Under local dependence, the dependence–adjusted FDP estimates for GAIF and Adaptive GAIF are, respectively:
$\widehat{\operatorname{FDP}}_{\mathrm{GAIF}_{\mathrm{dep}}}(t)$ $=\frac{\sum_{j\in\mathcal{I}_{t}}(1-\theta_{j})\alpha_{j}+\sum_{j\in\bar{\mathcal{I}}_{t}}\alpha_{j}}{\bigl(\sum_{j\leq t,\;j\notin\{t-L_{t},\dots,t-1\}}\delta_{j}\bigr)\vee 1},$
and
$\widehat{\operatorname{FDP}}_{\mathrm{Ada\text{-}GAIF}_{\mathrm{dep}}}(t)=\frac{\sum_{j<t-L_{t}}\frac{\alpha_{j}\mathbb{I}\{p_{j}>\lambda\}}{1-\lambda}\mathcal{W}_{j}+\sum_{t-L_{t}\leq j\leq t}\frac{\alpha_{j}}{1-\lambda}\mathcal{W}_{j}}{\bigl(\sum_{j\leq t,\;j\notin\{t-L_{t},\dots,t-1\}}\delta_{j}\bigr)\vee 1},$
where $\mathcal{W}_{j}=(1-\theta_{j})\mathbb{I}\{j\in\mathcal{I}_{t}\}+\mathbb{I}\{j\in\bar{\mathcal{I}}_{t}\}$. The corresponding testing levels are chosen so that $\widehat{\mathrm{FDP}}_{\mathrm{GAIF}_{\mathrm{dep}}}(t)\leq\alpha$ or $\widehat{\mathrm{FDP}}_{\mathrm{Ada\text{-}GAIF}_{\mathrm{dep}}}(t)\leq\alpha$. The resulting testing levels are provided in Appendix A.2. Theorem 3 shows that $\text{GAIF}_{\rm dep}$ and $\text{Ada-GAIF}_{\rm dep}$ methods control mFDR under local dependence.
@@
Suppose that the null $p$-values are locally dependent as defined in (11). Then if the test levels $\{\alpha_{t}\}_{t\in\mathbb{N}}$ satisfy $\widehat{\operatorname{FDP}}_{\operatorname{GAIF}_{\rm{dep}}}(t)\leq\alpha$ or $\widehat{\operatorname{FDP}}_{\operatorname{Ada-GAIF}_{\rm{dep}}}(t)\leq\alpha$, we have $\operatorname{mFDR}(t)\leq\alpha$ for all $t\in\mathbb{N}$.
4 Applications of GAIF on Online Conformal Selection
Online conformal selection [Huo et al., 2024, Xu and Ramdas, 2024] provides a canonical and intuitive instantiation of the GAIF framework. Here, decisions must be made in real time about whether an incoming observation satisfies a pre-specified requirement, while the true label is typically observed after the decision and thus serves as feedback. Throughout this section, we focus on the full and instant feedback setting.
Consider a data pair $({\bf X},Y)\in{\mathcal{X}}\times{\mathcal{Y}}$, with a historical calibration dataset $\mathcal{D}_{{\mathcal{C}}}=\{({\bf X}_{i},Y_{i})\}_{i=-n+1}^{0}$ of size $n$, with index set $\mathcal{C}$}. Test samples $({\bf X}_{t},Y_{t})$ arrive sequentially for $t=1,2,\dots$. At each time $t$, the covariate ${\bf X}_{t}$ is observed but responses $Y_{t}$ remain hidden until a real-time decision $\delta_{t}\in\{0,1\}$ is made. The goal is to determine whether $Y_{t}$ lies in a target region ${\mathcal{A}}\subseteq{\mathcal{Y}}$, for example, ${\mathcal{A}}=[a,b]$ or ${\mathcal{A}}=[b,\infty)$ in regression tasks. This can be framed as an online testing problem with $\theta_{t}=\mathbb{I}\{Y_{t}\in{\mathcal{A}}\}$.
We apply the GAIF framework to online conformal selection as follows. Since no off-the-shelf $p$-values are available, we first construct feedback-assisted online conformal $p$-values in Section 4.1. We then incorporate them into a modified GAIF (or Adaptive GAIF) rule to obtain finite-sample mFDR control in Section 4.2. We further use feedback to adaptively select the score model for conformal testing in Section 4.3, and analyze its optimality in Section 4.4.
@@ 4.1 Construction of online conformal $p$-values
We begin by constructing suitable online $p$-values for this online conformal selection setting. Let $V({\bf X})$ be a non-conformity score, where larger values indicate a higher likelihood that $\theta_{i}=0$. Typically, $V({\bf X})$ is a monotone transformation of the prediction $\widehat{\mu}({\bf X})$, assumed pre-trained to estimate $Y_{t}$. For example, if $\theta=\mathbb{I}\{Y_{t}\geq b\}$, one can take $V(X)=b-\widehat{\mu}({\bf X})$. For simplicity, we write $V_{i}=V({\bf X}_{i})$.
A natural approach to achieving online FDR control is to compute conformal $p$-values with the given calibration set $\mathcal{C}$} [Bates et al., 2023, Jin and Candès, 2023] and then apply GAIF or existing GAI rules. However, standard conformal $p$-values are not suitable here, since they do not satisfy the conditional super-uniformity in (6) and independence assumptions. To circumvent these issues, we adopt online conformal $p$-values, which are constructed by sequentially updating the calibration set. Specifically, at time $t$, the current point $({\bf X}_{t},Y_{t})$ serves as the test point, but from time $t+1$ onward, it is added to the calibration set. Let ${\mathcal{C}}_{0t}$ denote the dynamically updated index set of calibration samples with
${\mathcal{C}}_{0t}=\{-n<i<t:\theta_{i}=0\}.$
where $\xi_{t}\overset{\text{i.i.d.}}{\sim}\operatorname{Unif}[0,1]$ is an independent random variable for tie-breaking. Similar forms of online conformal $p$-values have been studied for testing exchangeability [Vovk et al., 2003, Vovk, 2021] and constructing online conformal prediction intervals [Angelopoulos et al., 2024].
Under exchangeability, this construction yields null $p$-values that are independent, thereby avoiding the complex dependence induced by structures inherent in offline conformal methods, which rely on shared calibration sets and produce $p$-values that are PRDS [Bates et al., 2023]. We impose the following assumption.
@@
This assumption is common in conformal inference [Marandon et al., 2024], and is weaker than requiring that the full data $\{({\bf X}_{i},Y_{i})\}_{i>-n}$ are independent and identically distributed (i.i.d.). We formally state the validity and independence under the null in Proposition 4.1.
@@
The mutual independence above indeed relies on the full and instant feedback setting, since only in this case does ${\mathcal{C}}_{0t}$ contain the complete set of past null samples, allowing the sequential rank-based independence argument to hold. For delayed feedback, additional adjustments are needed; see the finite-sample analysis in Appendix C.1. We also allow the score function to vary over time and be data-dependent, provided that it remains symmetric on the null samples; see Appendix F.5. This flexibility enables leveraging non-null feedback to improve score quality, facilitating score selection and adaptation to distribution shift.
4.2 Online conformal testing with feedback
Having established the properties of online conformal $p$-values, we now introduce the corresponding testing procedures. Although applying GAIF directly to those $p$-values yields satisfactory FDR control empirically, a finite-sample theoretical guarantee remains lacking. The key challenge is that null online conformal $p$-values may still be dependent on non-null decisions through the adaptive testing levels.
To overcome this issue, we propose a modified GAIF rule tailored for online conformal testing, offering rigorous, distribution-free, and model-agnostic mFDR guarantees. The key idea is to construct the sequence $\{\alpha_{t}\}_{t\in\mathbb{N}}$ based solely on the rejections of true nulls, rather than on all rejections. While this strategy may be conservative as discarding non-null decision information, the integration of the feedback provides substantial compensation.
More specifically, let $\tau_{j}$ be the time of the $j$-th rejection, and denote $\tilde{\tau}_{j}$ as the time of the $j$-th rejection under the null, defined as
@@ 4.2 Online conformal testing with feedback
We obtain the safe versions of LF and SF by replacing $\tau_{j}$ with $\tilde{\tau}_{j}$, leading to the LFS and SFS procedures, respectively. The letter “S" stands for “safe", since it guarantees rigorous finite-sample mFDR control. By doing so, $\tilde{\tau}_{j}$ is fixed given past null p-values, as its construction depends only on the rejections of true null hypotheses. The Online Conformal Testing with Feedback (OCTF) procedureis detailed in Algorithm 2.
@@
1:  Initialize ${\mathcal{C}}_{01}=\{i\in{\mathcal{C}}:\theta_{i}=0\}$.
@@
3:   Observe test covariate ${\bf X}_{t}$.
4:   Compute conformity scores $\{V_{i}\}_{i\in{\mathcal{C}}_{0t}\cup\{t\}}$.
5:   Compute online conformal $p$-value $p_{t}$ via (12).
6:   Update $\alpha_{t}={\alpha}_{t}^{\text{LFS}}$ in (18) (or $\alpha_{t}={\alpha}_{t}^{\text{SFS}}$ in (19)).
7:   if $p_{t}\leq\alpha_{t}$ then $\delta_{t}=1$, else $\delta_{t}=0$.
8:   Obtain the revealed feedback $Y_{t}$, and thus obtain $\theta_{t}$.
@@
Theorem 4 (Finite-sample online mFDR control for OCTF).
Suppose Assumption 1 holds. The output $\mathcal{R}$} of Algorithm 2 satisfies $\operatorname{mFDR}(t)\leq\alpha$ for all $t\in\mathbb{N}$.
With respect to FDR control, the additional challenge lies in the fact that a past non-null decision $\delta_{i}=\mathbb{I}\{p_{i}\leq\alpha_{i}\}$ with $\theta_{i}=1$ may be dependent on a current null $p_{t}$. Decoupling the dependence between decisions $\sum_{i\leq t}\delta_{i}$ in the denominator and $p_{t}$ is more challenging and warrants further research.
4.3 Optimized online conformal testing with feedback
The performance of online conformal testing depends on the choice of the non-conformity scores $V({\bf X})$ and the prediction model, e.g., random forests, neural networks, support vector machines, or regularized linear models [Bai and Jin, 2024, Gasparin and Ramdas, 2024]. In practice, the most effective score may also change over time, especially when the non-null distribution drifts [Gang et al., 2023, Huo et al., 2024]. This motivates a feedback-driven score-selection step built on top of the OCTF framework.
@@ 4.3 Optimized online conformal testing with feedback
$\widehat{k}_{t}=\underset{k\in[K]}{\arg\min}\;{\mathcal{M}}({\mathcal{D}}_{t};k),$
where ${\mathcal{M}}({\mathcal{D}}_{t};k)$ is a given criterion evaluated by the currently observed data ${\mathcal{D}}_{t}=\{({\bf X}_{i},Y_{i}):-n<i<t\}\cup\{{\bf X}_{t}\}$. Our goal is to choose the score that is most favorable for detecting non-nulls, which motivates the design of a reliable criterion ${\mathcal{M}}({\mathcal{D}}_{t};k)$. Since the testing threshold $\alpha_{t}$ depends recursively on the past rejection path, directly optimizing rejection probability is difficult. We therefore use the proxy
${\mathbb{E}}[p_{k,t}\mid\theta_{t}=1],$
where $p_{k,t}$ is the $t$-th conformal $p$-value obtained from the score function $V(\cdot;k)$. Smaller non-null $p$-values generally correspond to higher rejection probability $\Pr(p_{k,t}\leq\alpha_{t}\mid\theta_{t}=1)$, so this provides a natural threshold-free criterion.
To estimate ${\mathbb{E}}[p_{k,t}\mid\theta_{t}=1]$ under shifted non-null distribution while preserving symmetry, we employ an exponentially weighted moving average (EWMA) of past auxiliary non-null $p$-values. Specifically, at time $t$, for each score function $k$, we construct auxiliary non-null $p$-values as
$\tilde{p}_{k,j}=\frac{\sum_{s\in{\mathcal{C}}_{0t}\cup\{t\}}\mathbb{I}\{V(\mathbf{X}_{s};k)\leq V(\mathbf{X}_{j};k)\}}{1+|{\mathcal{C}}_{0t}|},\quad j\in{\mathcal{C}}_{1t},$ (13)
@@
where ${\mathcal{C}}_{1t}=\{-n<i<t:\theta_{i}=1\}$ denotes the set of online non-null samples observed prior to time $t$. When the non-null distribution varies slowly over time, these auxiliary non-null $p$-values are approximately distributed like the current non-null $p$-value $p_{k,t}$. The construction in (13) is carefully designed so that the resulting auxiliary non-null $p$-values are invariant under permutations of ${\mathcal{C}}_{0t}\cup\{t\}$, which is crucial for ensuring valid inference after score selection. We then define the EWMA criterion as
${\mathcal{M}}^{\mathrm{EWMA}}_{t}({\mathcal{D}}_{t};k)=\frac{\sum_{j\in{\mathcal{C}}_{1t}}\rho^{t-1-j}\,\tilde{p}_{k,j}\,}{\sum_{j\in{\mathcal{C}}_{1t}}\rho^{t-1-j}},$
@@
where $\rho\in(0,1)$ is the user-specified decay parameter that downweights past observations, enabling dynamic adaptation to the recent non‑null distribution and precise estimation of ${\mathbb{E}}[p_{k,t}\mid\theta_{t}=1]$. The selected score function is accordingly as $\widehat{k}_{t}={\arg\min}_{k\in[K]}{\mathcal{M}}^{\mathrm{EWMA}}_{t}({\mathcal{D}}_{t};k).$
The online conformal $p$-value after score selection, denoted as $p^{\rm opt}_{t}$, is then computed by (12) using the selected score $V(\mathbf{X}_{t};\widehat{k}_{t})$. Then running the OCTF procedures in Section 4.2 with these optimized $p$-values yields the whole optimized procedure, which we refer to as Opt-OCTF algorithm.